OpenStudy (anonymous):

Suppose 2 kg of a radioactive substance decays to 250 g in 75 years. What is the half–life of the substance?

OpenStudy (anonymous):

\[A_0=2000g\]\[A_1=250g\]so \[A_1=A_0 \times e^{kt}\]\[250=2000 \times e^{75k}\]\[\frac{250}{2000}=e^{75k}\]\[\ln(0.125)=75k\]\[\frac{\ln(0.125)}{75}=k \approx-0.028\]

OpenStudy (anonymous):

Thank you so much! :)